# Solvent: liquidity verification of smart contracts

### v1

@@ Abstract
Smart contracts are programs executed by blockchains networks to regulate the exchange of crypto-assets between untrusted users. Due to their immutability, public accessibility and high value at stake, smart contracts are an attractive target for attackers, as evidenced by a long history of security incidents. This has been a driving factor for the application of formal methods to Ethereum, the leading smart contract platform, and Solidity, its main smart contract language, which have become the target of dozens of verification tools with varying objectives. A current limitation of these tools is that they are not really effective in expressing and verifying liquidity properties regarding the exchange of crypto-assets: for example, is it true that in every reachable state a user can fire a sequence of transactions to withdraw a given amount of crypto-assets? We propose Solvent, a tool aimed at verifying these kinds of properties, which are beyond the reach of existing verification tools for Solidity. We evaluate the effectiveness and performance of Solvent through a common benchmark of smart contracts.
@@ 1 Introduction
In recent years we have seen a steady rise of decentralized applications that implement financial ecosystems on top of public blockchains. These applications are based on so-called smart contracts, i.e. programs that are run by blockchain networks and automatize the exchange of crypto-assets between users, without relying on trusted authorities. It is estimated that more than 100 billions of dollars worth of crypto-assets are today controlled by smart contracts [14]. The peculiarities of the setting (i.e., the absence of intermediaries, the immutability of code after deployment, the quirks in smart contract languages) make smart contracts an appealing target for attackers, as bugs might be exploited to steal crypto-assets or just cause disruption. This is witnessed by a long history of attacks, which overall caused losses exceeding 6 billions of dollars [12].
@@
Formal methods provide an ideal defense against these attacks, since they enable the creation of tools to detect bugs in smart contracts before they are deployed. Indeed, smart contracts verification tools, often based on formal methods, have been mushrooming in the last few years: for Ethereum alone — largely the main smart contract platform — dozens of tools exist today [20]. Still, the actual effectiveness of these tools in countering real-world attacks is debatable: indeed, attacks to smart contracts have continued to proliferate, refining their strategies from exploits of known vulnerability patterns to sophisticated attacks to the contracts’ business logics. As a matter of fact, the vast majority of the losses due to real-world attacks are caused by logic errors in the contract code [12], which are outside the scope of most vulnerability detection tools.
The main tools that support the verification of the contracts’ business logic have serious limitations concerning soundness, completeness and expressivity [8]. The reasons of these limitations are in part inherent to the target languages of the tools (Solidity and its compiled bytecode), which have some design choices that are problematic for verification. The well-known reentrancy issue, which is an ongoing cause of problems since the infamous DAO attack [5], is only one of the causes: in general, the glitches of high-level abstractions over the low-level target (e.g., gas costs [15], vulnerable tokens [28]) are a common cause of bugs. Second, existing verification tools for Solidity have limited support for properties asking that, whenever certain states are reached, users can perform some actions that lead to a desirable asset exchange [24]. Properties of this kind, which are neither safety nor liveness properties and are often dubbed liquidity [26, 9, 21], are crucial in the smart contracts setting: typically, one would like to know if, after having deposited some crypto-assets in a contract, they will be able to redeem a given function of the deposited amount. Not being able to verify these properties is a serious limitation: indeed, several real-world attacks exploited liquidity weaknesses, allowing attackers to steal or freeze crypto-assets [3].
To address these issues we propose Solvent  [^1] https://github.com/AngeloFerrando/Solvent, a tool that verifies liquidity properties of smart contracts: in particular, Solvent can practically express and verify properties of the exchanges of assets between users. To overcome the above-mentioned limitations due to Solidity, Solvent operates on a version of Solidity purified from the main semantical quirks of the full language, and extended with special constructs to express hashed timelock protocols, which are commonly used in contracts involving user-chosen secrets (e.g., in gambling games and blind auctions). Solvent takes as input a contract and a set of user-defined liquidity properties, and translates them into constraints in the SMT-LIB standard [7], reducing the verification problem to a symbolic model checking one. Then, techniques such as bounded model checking and predicate abstraction are employed, relying on Z3 [13] and cvc5 [6] as back-end SMT solvers. Experiments on a benchmark of real-world smart contracts show that Solvent can efficiently verify relevant liquidity properties of their behaviour. These properties are out of the scope of industrial verification tools operating on the full Solidity, like e.g. SolCMC [4] and Certora [18], as well as academic tools [19, 16, 23, 22, 25, 27] (due to space constraints, we compare Solvent with these tools in the Appendix). Solvent provides developers with useful feedback, by detecting logical errors that would otherwise remain unnoticed. In particular, when Solvent detects that a property is violated, it produces a witness of the violation in terms of a concrete execution trace leading the contract to a state from which the desired state change (possibly involving transfers of crypto-assets) is not realisable.
@@
The main contributions of this tool demostration paper include:
@@
2 Tool demo
In this section we demonstrate our tool through a simple example and provide some highlights on how it works. Solvent operates in two steps: 1 given as input a smart contract and a set of liquidity properties, it encodes the contract and the properties into SMT constraints; 2 then, it issues satisfiability queries to SMT solvers (in our setting, Z3 and cvc5) to detect if the required properties are violated. If so, it produces a counterexample, in the form of a sequence of transactions leading to a state where a required property cannot be satisfied.
@@ 2 Tool demo
Forall xa [
!target_reached && block.number>end_donate
-> Exists tx [1, xa]
[ <tx>balance[xa] >= balance[xa]+donors[xa] ] ]
@@
To illustrate Solvent, we consider in LABEL:lst:crowdfund-bug a simple crowdfunding contract. The contract is akin to a class in OO programming, with attributes that define its state and methods (triggered by blockchain transactions) that update it. The user who fires the transaction (denoted by msg.sender) can transfer some ETH to the contract along with the call (denoted by msg.value). The constructor specifies the owner of the crowdfunding campaign, a deadline for donations, and the target (in ETH, the Ethereum native crytocurrency). The method donate allows anyone to donate any ETH amount before the deadline; wdOwner allows the owner to redeem the whole contract balance if the campaign target has been reached and the deadline has expired; finally, wdDonor allows donors to withdraw their donations after the deadline, if the campaign target has not been reached.
@@
To translate donor_wd into an SMT constraint, Solvent considers its negation, and introduces a new existentially quantified variable for xa, and new universally quantified variables for all transaction variables in the sequence tx and for all next-state variables. Then, it reduces to checking whether there exists an xa for which, if the antecedent holds, for all transaction variables and next-state variables, either the transactions invalidly tie current and next-state, or the required consequent does not hold. If this formula is unsatisfiable, then the property holds; otherwise, a counterexample is produced. E.g., for donors_wd, the counterexample is the following sequence of transactions, which lead to a state where the property cannot be satisfied:
setfontsize8.59.5
[1] constructor(0,2) msg.sender=address(4) msg.value=0
@@
[3] coinbase() msg.sender=address(0) msg.value=1
@@ 3 Evaluation
We test our tool over a common benchmark for Solidity verification [2], which includes a representative set of real-world contracts and properties. Since this benchmark is focussed on current verification tools for Solidity, which have limited support for liquidity properties, we extend it with relevant properties of this kind for each contract (see the github page). Overall, we end up with 107 verification tasks, which we manually check for the ground truth.
@@ 4 Conclusions and Future work
Solvent has already proven useful to spot non-trivial liquidity vulnerabilities in smart contracts, which are beyond the reach of mainstream verification tools for Solidity. Still, there is space for improvements. First, Solvent makes use of off-the-shelf SMT solvers, relying on bounded model checking to find counterexamples, and on predicate abstraction to prove that the property holds. An alternative approach is to leverage more advanced techniques (such as abstraction-refinement, k-induction, etc.) used by modern infinite-state symbolic model checkers. Further future work involves extending the Solvent fragment to narrow the gap with the full Solidity language. This could amount to extend the subset of Solidity supported by the tool to include e.g. contract-to-contract calls (see the Appendix for further extensions), and to automatically transform the transaction sequences outputted by Solvent into executable Ethereum tests in the Ganache tool.
@@ Appendix 0.A Supplementary material
  • in Section 0.A.2 we discuss related work on the verification of smart contracts. In particular, we draw a detailed comparison between Solvent and the two main verification tools for Solidity, i.e. SolCMC and Certora, highlighting their weaknesses in the verification of liquidity properties. We also compare Solvent with other verification tools that can address liveness properties, such as VeriSolid [22] and SmartPulse [25].
  • in Section 0.A.3 we evaluate Solvent on a more complex use case, i.e. a fair 2-players lottery. Implementing this use case is quite error-prone, since the contract must implement a commit-reveal protocol that prescribes punishments whenever a player behaves dishonestly, e.g. by refusing to perform some required action. The contract must ensure that, even in these cases, an honest player has at least the same payoff that she would have by interacting with another honest player. Proving relevant liquidity properties of this contract seems far beyond the capabilities of existing verification tools. To overcome these limitations, we provide Solvent with an extended syntax to express hashed timelock protocols, similarly to what is done in other smart contract languages (e.g., in Tezos).
@@ 0.A.1 Background on Solidity, and relation to Solvent’s fragment
  • In Solidity, transfers of crypto-currency from a contract to another account are encoded as contract-to-contract calls, while in Solvent we assume that the transferred amount actually arrives at the destination address, i.e. that this address is either an EOA or a contract account that does not perform a further call. This assumption simplifies stating properties about the funds transferred from a contract (see Section 0.A.2), and is used in other verification tools such as VerX [23], where it is called effective external callback freedom assumption.
@@
  • Solvent features a special syntax to express hashed timelock protocols, where a user first store the hash of a chosen secret in the contract, and then reveal the secret, making the contract check that the hash of the revealed secret corresponds to the stored hash. This is a common pattern, used e.g. in gambling games and blind auctions, but where existing verification tools for Solidity fail to give the expected results (see Section 0.A.2). We exemplify this feature to design a 2-players lottery in Section 0.A.3, showing that Solvent manages to provide developers with a useful feedback.
0.A.2 Related work
@@
Current mainstream verification tools for Solidity are quite limited when it comes to liquidity properties. SolCMC, the prover shipped with the Solidity compiler, and other verification tools such as Zeus [19], solc-verify [16], VerX [23] and SmartACE [27], support safety properties, and therefore they cannot express any of the liquidity properties considered in this paper.
Other tools such as Certora [18], VeriSolid [22], and SmartPulse [25], can reason about liveness properties, i.e. properties of the form “for all reachable states, eventually some desirable outcome will be obtained”. While liveness properties are closely related to liquidity properties, and in some cases can be used to express similar concepts, they cannot express most of the properties considered in this paper (and relevant in practice): namely, they cannot express properties of the form “for all reachable states, some user can do something that immediately produces some desirable effect”.
@@
First, we analyse Certora, one of the leading formal verification tools for Solidity. To illustrate the expressiveness limitations of the tool, consider the Freezable contract in LABEL:lst:freezable-deposit. The contract allows anyone to withdraw part of its balance through the method pay, unless the variable frozen is true. This variable is controlled by the owner through the method freeze: hence, the owner at any time can freeze the contract balance, preventing anyone from withdrawing. A desirable property of the Freezable contract, and of smart contracts in general, is that crypto-assets cannot be frozen forever.
@@
[ <tx> balance[xa] == balance[xa] + balance ]
@@
<tx>balance[rcv] == balance[rcv] + balance
@@
Besides Certora, only SmartPulse [25] and VeriSolid [22] seem capable to deal with liveness properties, which are similar (though not equivalent) to liquidity.
SmartPulse targets directly Solidity code, and it has a property specification language based on Linear Temporal Logic (LTL). This makes it possible to express liveness properties, but not liquidity properties, that are out of the reach of LTL. Indeed, in order to express the concept “there exists a path that leads to a desirable outcome”, one needs to use the E temporal operator of CTL. Therefore, the properties supported by SmartPulse and Solvent have uncomparable expressiveness. To be usable in practice, liveness properties must be accompanied by a fairness assumption, i.e. another LTL formula that specifies the traces where a user performs some required action in order to reach the desired state. For example, in our crowdfunding contract (Section 2) a target property could be “if the target is not reached, then a donor gets its money back”, and the associated fairness assumption could be “the donor performs the action wdDonor”. A main difference between Solvent and SmartPulse is that, while in SmartPulse the designer of the property must anticipate, in the fairness assumption, the sequence of transactions that a user must perform in order to reach a certain state change, in Solvent this sequence is inferred by the the SMT solver. In particular, the solver infers the (minimal) length of the sequence, the called methods, and their actual parameters in order to produce the desired state change. This allows for a more succinct representation of liveness properties compared to SmartPulse.
VeriSolid takes as input a Solidity contract and its properties expressed in Computation Tree Logic (CTL), transforms the contract into an equivalent Abstract State Machine (ASM), and verifies the properties against the ASM using tools in the BIP toolchain, such as the nuXmv symbolic model checker [10]. The liveness properties specified in [22] are not accompanied by fairness assumptions (unlike SmartPulse), but in principle this seems doable without reworking the verification techniques. [^5] A recent extension of VeriSolid include some forms of fairness constraints [11]. It is not clear whether the liveness properties expressed in CTL can encompass the liquidity properties addressed by Solvent: indeed, to express liquidity we mix universal and existential quantification on variables (e.g., “for all users, there exists a sequence of transactions”), which seems beyond the scope of plain CTL.
0.A.3 Use case: a 2-players lottery

### v2

@@ Abstract
Smart contracts are an attractive target for attackers, as evidenced by a long history of security incidents. A current limitation of smart contract verification tools is that they are not really effective in expressing and verifying liquidity properties regarding the exchange of crypto-assets: for example, is it true that in every reachable state a user can fire a sequence of transactions to withdraw a given amount of crypto-assets? We propose Solvent, a tool aimed at verifying these kinds of properties, which are beyond the reach of existing verification tools for Solidity. We evaluate the effectiveness and performance of Solvent through a common benchmark of smart contracts.
@@ 1 Introduction
In recent years we have seen a steady rise of smart contracts that implement financial ecosystems on top of public blockchains, and control today more than 100 billions of dollars worth of crypto-assets [15]. The peculiarities of the setting (i.e., the absence of intermediaries, the immutability of code after deployment, the quirks in smart contract languages) make smart contracts an appealing target for attackers, as bugs might be exploited to steal crypto-assets or just cause disruption. This is witnessed by a long history of attacks, which overall caused losses exceeding 6 billions of dollars [13].
@@
Formal methods provide an ideal defense against these attacks, since they enable the creation of tools to detect bugs in smart contracts before they are deployed. Indeed, smart contracts verification tools, often based on formal methods, have been mushrooming in the last few years: for Ethereum alone — largely the main smart contract platform — dozens of tools exist today [23]. Still, the actual effectiveness of these tools in countering real-world attacks is debatable: indeed, attacks to smart contracts have continued to proliferate, refining their strategies from exploits of known vulnerability patterns to sophisticated attacks to the contracts’ business logics. As a matter of fact, the vast majority of the losses due to real-world attacks are caused by logic errors in the contract code [13], which are outside the scope of most vulnerability detection tools. In particular, several real-world attacks were based on liquidity weaknesses of smart contracts, which were exploited by attackers to steal or freeze crypto-assets [3]. Liquidity expresses the ideal behaviour of contracts in terms of the exchange of crypto-assets [34, 10, 24]: users want to be guaranteed that, whenever certain states are reached, they can always perform some actions that lead to a desirable assets transfer [29]. The resulting class of properties is dubbed, in general, liquidity.
The current tools that support the verification of behavioural properties of Solidity contracts cannot verify general liquidity properties. This is due both to the design choices of Solidity, their target language, and to the complex logical structure of such properties. The difficulty of verifying Solidity is caused by several glitches of high-level abstractions over the low-level target (e.g. reentrancy [5], gas costs [18], and non-native tokens [36]). Indeed, it has been observed that existing tools already face challenges (in terms of soundness and completeness) in the verification of even simpler properties [9]. We believe that effectively verifying liquidity requires first to purify Solidity from its main semantical quirks. Once we have done this, we must deal with the peculiar logical structure of liquidity properties. We start by noting that liquidity cannot be expressed in terms of safety or liveness (see also Section 0.A.2). Indeed, while safety has the form “$p$ always holds”, and liveness has the form “$p$ eventually holds”, liquidity instead has the form “for certain users $a$, there always exists a sequence of at most $m$ actions that $a$ can perform to make $p$ hold” (where $m$ is a parameter). Notice that bounded liveness cannot model liquidity either: not being able to branch over paths, it can only express properties of the form “$p$ always holds after at most $m$ steps”. Indeed, this may even be undesirable: for example, a bank client does not want to “always withdraw their deposit after at most $m$ steps” (maybe they want to keep their money deposited for $m+1$ steps); rather, they want to know whether they “can always withdraw their deposit, by performing at most $m$ actions”. To deal with liquidity, verification techniques that go beyond safety and (bounded) liveness are therefore necessary.
@@
We propose Solvent  [^1] https://github.com/AngeloFerrando/Solvent, a tool that verifies liquidity properties of smart contracts. Solvent takes as input a contract, written in a purified version of Solidity, and a set of user-defined liquidity properties. The tool translates them into SMT constraints [8], reducing the verification problem to a symbolic model checking one. Then, techniques such as bounded model checking and predicate abstraction are employed, relying on Z3 [14] and cvc5 [6] as back-end SMT solvers. Experiments on a benchmark of real-world smart contracts show that Solvent can efficiently verify relevant liquidity properties of their behaviour. These properties are currently out of the scope of industrial verification tools operating on the full Solidity, like e.g. SolCMC [4] and Certora [21], as well as academic tools [22, 19, 28, 25, 31, 35] (due to space constraints, we discuss such tools in Section 0.A.3). Solvent provides developers with useful feedback, by detecting logical errors that would otherwise remain unnoticed. In particular, when Solvent detects a property violation, it produces a concrete execution trace that leads the contract to a state from which the desired asset exchange is unrealisable.
Summing up, the main contributions of the paper include:
@@
2 Verifying liquidity properties with Solvent
We demonstrate our tool through a simple example and provide some highlights on how it works. Solvent operates in two steps: 1 given as input a smart contract and a set of liquidity properties, it encodes the contract and the properties into constraints in the SMT-LIB standard [7]; 2 then, it issues satisfiability queries to an SMT solver to detect if the required properties are violated. If so, it produces a counterexample, in the form of a sequence of transactions leading to a state where a required property cannot be satisfied.
To illustrate Solvent, we consider in LABEL:lst:crowdfund-bug a simple crowdfunding contract. The contract is akin to a class in OO programming, with attributes that define its state and methods (triggered by blockchain transactions) that update it. The user who fires the transaction (denoted by msg.sender) can transfer some amount (msg.value) of cryptocurrency to the contract along with the call. The constructor specifies the owner of the crowdfunding campaign, the deadline for donations, and the target amount. The method donate allows anyone to donate any amount before the deadline; wdOwner allows the owner to redeem the whole contract balance if the campaign target has been reached and the deadline has expired; finally, wdDonor allows donors to withdraw their donations after the deadline, if the campaign target has not been reached.
@@ 2 Verifying liquidity properties with Solvent
Forall xa [ !target_reached && block.number>end_donate
-> Exists tx [1,xa] // 1 = max tx length, xa = tx sender
[ <tx>xa.balance // balance of xa after tx is performed
>= xa.balance + donors[xa] ] ]
@@
To translate donor_wd into an SMT constraint, Solvent considers its negation, and introduces a new existentially quantified variable for xa, and new universally quantified variables for all transaction variables in the sequence tx and for all next-state variables. Then, it reduces to checking if there exists an xa for which, if the antecedent holds, for all transaction variables and next-state variables, either the transactions invalidly tie current and next-state, or the required consequent does not hold. If this formula is unsatisfiable, then the property holds; otherwise, a counterexample is produced. E.g., for donors_wd, the counterexample is the following sequence of transitions. [^2] Transition 3 represents a call to an adversarial contract, with a method invoking selfdestruct on the crowdfund contract. We note that selfdestruct is still active [16]. This can be easily translated into a concrete PoC, leading the contract to a state where the property cannot be satisfied.
setfontsize89
[1] constructor(2,0,2) msg.sender=address(4) msg.value=0
@@
[3] selfdestruct() msg.sender=address(0) msg.value=1
@@ 3 Evaluation
We test our tool over a common benchmark for Solidity verification [2], which includes a representative set of real-world contracts and properties. Since this benchmark is focused on current verification tools for Solidity, which do not deal with general liquidity properties, we extend it with relevant properties of this kind for each contract (see the github page). Overall, we end up with 107 verification tasks, which we manually check for the ground truth.
@@ 4 Conclusions and Future work
Solvent has already proven useful to spot general liquidity vulnerabilities, which are beyond the reach of current verification tools. Still, there is space for improvements. First, Solvent uses off-the-shelf SMT solvers, relying on bounded model checking to find counterexamples, and on predicate abstraction to prove that the property holds. An alternative approach is to leverage advanced techniques (e.g., abstraction-refinement, k-induction, etc.) used by modern infinite-state symbolic model checkers. Further work involves extending the Solvent fragment to narrow the gap with the actual Solidity. This could amount to extend the supported Solidity fragment to include e.g. contract-to-contract calls (see Appendix A for further extensions), and to automatically transform the transaction sequences outputted by Solvent into an executable PoC in the HardHat tool. The formalization of liquidity properties in a temporal logic is left as future work.
@@ Appendix 0.A Supplementary material
  • in Section 0.A.2 we discuss why liquidity properties cannot be expressed as safety or liveness properties
  • in Section 0.A.3 we discuss related work on bug detection and formal verification tools for smart contracts. In particular, we provide a thoroughly discussion of the differences between Solvent and the two main verification tools for Solidity, i.e. SolCMC and Certora, highlighting why verification of liquidity properties is currently out of their scope. We also discuss other verification tools, such as VeriSolid [25] and SmartPulse [31], that can verify liveness properties but not liquidity ones.
  • in Section 0.A.4 we evaluate Solvent on a more complex use case, i.e. a fair 2-players lottery. Implementing this use case is quite error-prone, since the contract must implement a commit-reveal protocol that prescribes punishments whenever a player behaves dishonestly, e.g. by refusing to perform some required action. The contract must ensure that, even in these cases, an honest player has at least the same payoff that she would have by interacting with another honest player. Proving relevant liquidity properties of this contract seems far beyond the capabilities of existing verification tools. To overcome these limitations, we provide Solvent with an extended syntax to express hashed timelock protocols, similarly to what is done in other smart contract languages (e.g., in Tezos).
@@ 0.A.1 Background on Solidity, and relation to Solvent’s fragment
  • In Solidity, transfers of crypto-currency from a contract to another account are encoded as contract-to-contract calls, while in Solvent we assume that the transferred amount actually arrives at the destination address, i.e. that this address is either an EOA or a contract account that does not perform a further call. This assumption simplifies stating properties about the funds transferred from a contract (see Section 0.A.3), and is used in other verification tools such as VerX [28], where it is called effective external callback freedom assumption.
@@
  • Solvent features a special syntax to express hashed timelock protocols, where a user first store the hash of a chosen secret in the contract, and then reveal the secret, making the contract check that the hash of the revealed secret corresponds to the stored hash. This is a common pattern, used e.g. in gambling games and blind auctions, but where existing verification tools for Solidity fail to give the expected results (see Section 0.A.3). We exemplify this feature to design a 2-players lottery in Section 0.A.4, showing that Solvent manages to provide developers with a useful feedback.
0.A.2 Liquidity vs. Safety and Liveness
In this section, we support our claim that liquidity properties cannot be expressed in terms of safety or liveness. First, we provide an intuitive motivating example; then, we discuss the structural differences between liquidity and safety and liveness properties in terms of LTL and CTL.
Consider a user opening a bank account. The user wants to be guaranteed that, if they deposit a certain amount of money in the bank, then they will always be able to withdraw their money by performing a reasonable amount of actions, not that they always withdraw their money, nor that they will eventually withdraw their money. Requiring liveness but not liquidity would allow the bank to give a guarantee such as “any client will withdraw their money after the bank has been visited overall 1 million times after the deposit”, making the liveness property “eventually I will withdraw” true (assuming the fairness condition that visits to the bank happen infinitely often). But it does not make true the more reasonable and desirable liquidity property “I’m always able to perform at most 3 actions that make me withdraw”. Note that bounded liveness would not be an appropriate specification either. Indeed, the user does not want to “always withdraw after 3 actions are made” (the 3 actions could be “deposit more money”, or they could be made by other users). Branching is necessary, as we need to existentially quantify over the paths.
In terms of LTL, we have that safety properties have the form “${G}p$”, and liveness properties have the form “${GF}p$”. Liquidity properties, however, cannot be expressed in LTL, as we need to existentially quantify over the paths. In CTL, the closest formulation would be “${AG\ EX\ p\lor AG\ EX\ EX\ p\lor\dots\lor AG\ (EX)^{m}\ p}$” (with $m$ fixed). This formulation, however, still does not capture the strategic aspect of liquidity, as we require that the sequence that makes $p$ hold consists of actions made by a certain user, and therefore we need to impose conditions over transition variables.
0.A.3 Related work
The Ethereum ecosystem includes several bug detection tools that can spot specific forms of liquidity vulnerabilities in smart contracts. In that setting, these vulnerabilities are often referred to as “Locked Ether”, which are roughly described as the absence of a mechanism to withdraw Ether from the contract. The tools capable of detecting Locked Ether bugs include Slither [17], SmartCheck [32], Maian [27], Securify2 [34], ConFuzzius [33] and sFuzz [26]. Each of these tools has its own internal encoding of the Locked Ether property, making it difficult to compare them [30]. There are two main differences between these tools and ours. First, while these tools have an hard-coded Locked Ether property, Solvent can verify custom liquidity properties, defined by our specification language. Second, bug detection tools only focus on property violations, while Solvent (a formal verification tool) in some cases can successfully verify that a liquidity property is satisfied.
@@ 0.A.3 Related work
Current formal verification tools for Solidity do not support general liquidity properties. SolCMC, the prover shipped with the Solidity compiler, and other verification tools such as Zeus [22], solc-verify [19], VerX [28] and SmartACE [35], only support safety properties, and therefore they cannot express any of the liquidity properties considered in this paper.
Other tools such as Certora [21], VeriSolid [25], and SmartPulse [31], can also verify liveness properties. In the rest of this section we discuss more concretly why Certora, SmartPulse and VeriSolid cannot express general liquidity properties.
@@
First, we discuss Certora, one of the leading formal verification tools for Solidity. Consider the Freezable contract in LABEL:lst:freezable-deposit. The contract allows anyone to withdraw part of its balance through the method pay, unless the variable frozen is true. This variable is controlled by the owner through the method freeze: hence, the owner at any time can freeze the contract balance, preventing anyone from withdrawing. A desirable property of the Freezable contract, and of smart contracts in general, is that crypto-assets cannot be frozen forever.
@@
[ <tx>xa.balance == xa.balance + balance ]
@@
<tx>rcv.balance == rcv.balance + balance
@@
Besides Certora, only SmartPulse [31] and VeriSolid [25] seem capable to deal with liveness properties. SmartPulse targets directly Solidity code, and it has a property specification language based on Linear Temporal Logic (LTL). This makes it possible to express liveness properties, but not liquidity properties, that are out of the reach of LTL (see Section 0.A.2). Therefore, the properties supported by SmartPulse and Solvent have uncomparable expressiveness. To be usable in practice, liveness properties must be accompanied by a fairness assumption, i.e. another LTL formula that specifies the traces where a user performs some required action in order to reach the desired state. For example, in our crowdfunding contract (Section 2) a target property could be “if the target is not reached, then a donor gets its money back”, and the associated fairness assumption could be “the donor performs the action wdDonor”. A main difference between Solvent and SmartPulse is that, while in SmartPulse the designer of the property must anticipate, in the fairness assumption, the sequence of transactions that a user must perform in order to reach a certain state change, in Solvent this sequence is inferred by the the SMT solver. In particular, the solver infers the (minimal) length of the sequence, the called methods, and their actual parameters in order to produce the desired state change.
VeriSolid takes as input a Solidity contract and its properties expressed in Computation Tree Logic (CTL), transforms the contract into an equivalent Abstract State Machine (ASM), and verifies the properties against the ASM using tools in the BIP toolchain, such as the nuXmv symbolic model checker [11]. The liveness properties specified in [25] are not accompanied by fairness assumptions (unlike SmartPulse), but in principle this seems doable without reworking the verification techniques. [^6] A recent extension of VeriSolid include some forms of fairness constraints [12]. We have already noted that the liveness properties expressed in CTL cannot encompass the general liquidity properties addressed by Solvent. We further note that to express liquidity we mix universal and existential quantification on variables (e.g., “for all users, there exists a sequence of transactions made by the user”).
0.A.4 Use case: a 2-players lottery